There are two classroom A and B .if 10 students are sent from A to B, the number of students in each room become the same .if 20 students are sent from B to A , the number of students in A becomes double the number of students in B .Find the number of students in each room .
Answer by ankor@dixie-net.com(22740)

There are two classroom A and B .
if 10 students are sent from A to B, the number of students in each room become the same .
if 20 students are sent from B to A , the number of students in A becomes double the number of students in B .
Find the number of students in each room .
:
Let a = original no. of students in A
Let b = original no. of students in B
:
Write an equation for each statement, put eq in standard form
:
"There are two classroom A and B .if 10 students are sent from A to B, the number of students in each room become the same ."
a - 10 = b + 10
a - b = 10 + 10
a - b = 20
:
"If 20 students are sent from B to A , the number of students in A becomes double the number of students in B
A + 20 = 2(B-20)
A + 20 = 2b - 40
A - 2B = -40 - 20
A - 2B = -60
Subtract the above equation from the 1st equation
a - b = 20
a - 2b = -60
--------------subtacting eliminates a, find b
b = 80 students in B
and
a - 80 = 20
a = 20 + 80
a = 100 students in A
:
You should check the solutions in both statements